OpenStudy (krealsey):

Help me please! I will give you a medal and a fan. I only need help with C-F!

OpenStudy (krealsey):

In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x) Complete the table of values. x 100 + 23x 90(1.2x) 0 100+0 90(0) 1 100+23 90(1.2) 2 100+46 90(1.44) 3 100+69 90(1.728) 4 100+92 90(2.0736) 5 100+115 90(2.48832) B. Use the table to determine at approximately which point the number of cells will be the same for each organism. (I HAVE ALL ANSWERS ABOVE THIS) C. Graph the system of equations and show the point of intersection. D. Explain what the points graphed for each line represent. E. Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c. F. Find the point(s) of intersection, and explain what the intersection represents in the context of the problem.

OpenStudy (krealsey):

@ganeshie8 @hartnn

OpenStudy (krealsey):

@shansha1

OpenStudy (krealsey):

@Kainui

OpenStudy (krealsey):

@Nnesha @Qwertty123

OpenStudy (krealsey):

Someone please help

ganeshie8 (ganeshie8):

I think there is a mistake in your equation for \(m(x)\) Shouldn't it be \[\large 90(1.2)^x\] ?

OpenStudy (krealsey):

i don't think so because that is what I was given.

OpenStudy (krealsey):

I didn't come up with it, it was given to me

ganeshie8 (ganeshie8):

Could you take a screenshot of the original question and attach ?

OpenStudy (krealsey):

In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x) Complete the table of values. x 100 + 23x 90(1.2x) 0 1 2 3 4 5 Use the table to determine at approximately which point the number of cells will be the same for each organism. Graph the system of equations and show the point of intersection. Explain what the points graphed for each line represent. Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c. Find the point(s) of intersection, and explain what the intersection represents in the context of the problem.

OpenStudy (krealsey):

Thats the original question

ganeshie8 (ganeshie8):

I need you to take a screenshot and attach the pic here I think some key information is lost while copy pasting

OpenStudy (krealsey):

I just took it from the actual assignment. THIS IS ALL I AM GIVEN. My computer doesn't screenshot.

ganeshie8 (ganeshie8):

Then can you double check the m(x) equation

OpenStudy (krealsey):

It is the one I was given. I cannot submit the wrong question that was given to me, it is what it is.

ganeshie8 (ganeshie8):

I want you check this one in particular : ``` m(x) = 90(1.2x) ``` Is that `x` really at the bottom and not at the upper right of 1.2 ?

OpenStudy (krealsey):

It is in the upper right hand corner. But I cannot change this on my compter because it doesn't have the right tools to do so. The teacher knows this and knows what I mean.

OpenStudy (krealsey):

When I copy and paste it turns into what is above and I can't change it.

ganeshie8 (ganeshie8):

You cannot write 1.2x when you mean 1.2^x.

ganeshie8 (ganeshie8):

you may use the up arrow to write exponent

ganeshie8 (ganeshie8):

Thre is one mistake in your table too

ganeshie8 (ganeshie8):

when x = 0 : 90(1.2^0) = 90(1) = 90

ganeshie8 (ganeshie8):

you have got 0 incorrectly for the same

ganeshie8 (ganeshie8):

double check..

OpenStudy (krealsey):

I understand that, but when I copy it the arrow goes away. Lets move on please I know what to do for all the other questions except for the ones I asked for.

OpenStudy (krealsey):

I do see the mistakes, But i can't do anything about them on here, only in my document. BU tin my document they are right.

ganeshie8 (ganeshie8):

Okay, what is your answer for part B ?

OpenStudy (krealsey):

I already have that in my document. I only need C-F and help with one other small question.

ganeshie8 (ganeshie8):

I need to know the answer for part B before proceeding with Part c

OpenStudy (krealsey):

My computer is irritating me hold on

OpenStudy (krealsey):

It wont let me submit my answer to you

ganeshie8 (ganeshie8):

Nvm then lets look at part c

ganeshie8 (ganeshie8):

`C. Graph the system of equations and show the point of intersection.`

ganeshie8 (ganeshie8):

you may use below site to graph the equations : https://www.desmos.com/calculator

ganeshie8 (ganeshie8):

In that link enter below on the top left corner window : ``` y = 100 + 23x ```

ganeshie8 (ganeshie8):

for the second graph, enter ``` y = 90(1.2)^x ```

OpenStudy (krealsey):

Ok. I can only type short answers on here which is making me irritated. How do I submit the graph though or download it to submit?

ganeshie8 (ganeshie8):

Here it is : https://www.desmos.com/calculator/j0dgiia9xr

ganeshie8 (ganeshie8):

you will need to take a screenshot

OpenStudy (krealsey):

Okay. I might have trouble with that, but can you help me with the rest of the questions?

ganeshie8 (ganeshie8):

`D. Explain what the points graphed for each line represent.` Looks straightforward, have you tried answering this ?

ganeshie8 (ganeshie8):

The answer to that part is given in the question itself

ganeshie8 (ganeshie8):

` the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: `

ganeshie8 (ganeshie8):

Something like this : `the x coordinate of the points represent the week and the y coordinates represent the number of organisms in that week`

OpenStudy (krealsey):

Okay, what about E and F

OpenStudy (krealsey):

I also have one more question if you'd be willing to help after that. It is a small one.

OpenStudy (krealsey):

@ganeshie8

ganeshie8 (ganeshie8):

`E. Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c.`` The x coordinate of the intersection of the graphs is the solution to the equation 100 + 23x = 90(1.2)^x. The graphs are intersecting at x = 4.4. So the solution is 4.4

ganeshie8 (ganeshie8):

That covers both E and F actually

ganeshie8 (ganeshie8):

Just add this for part E : The graphs are intersecting at x = 4.4. So the solution is 4.4. This point represents the week in which both the organisms have same population.

OpenStudy (krealsey):

Okay. can you help me with one more question that is much quicker?

OpenStudy (krealsey):

A private school opened in 2005 with an initial enrollment of 85 students. The enrollment has increased by an average of 18 students each year since the school opened. A nearby public school had an enrollment of 95 in 2005, and its enrollment has increased by an average of 15 students per year. Let t = time in years, with t = 0 representing the year 2005. Let f(t) = the number of students enrolled at the private school and g(t) = the number of students enrolled at the public school. a. Create the two functions to represent the situation. b. Graph the system of equations to determine the the point of intersection. c. Interpret the point of intersection in the context of the problem. (I ONLY NEED HELP WITH C.)

ganeshie8 (ganeshie8):

The point of intersection determines the year in which both schools will have the same number of students

OpenStudy (krealsey):

Thank you so much.

ganeshie8 (ganeshie8):

yw

OpenStudy (krealsey):

@ganeshie8 For the graphing part of y=100+23 is there more to it? because it is just a flat line.